# Distribution-free data-driven smooth tests without $\chi^{2}$

### v2

@@ 4.1 On the construction of the K2 orthonormal basis
From Proposition 5, it follows that each function $\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$ fulfills (21), and thus we obtain
$\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\widetilde{\phi}_{j\bm{\gamma}}$ $=\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\left[\phi_{j\bm{\gamma}}-\bm{a}^{T}_{\bm{\gamma}}\left\langle\bm{a}^{T}_{\bm{\gamma}},\phi_{j\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}\right]$
$=\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}-\bm{b}^{T}_{\bm{\beta}}\left\langle\bm{b}^{T}_{\bm{\beta}},\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}}.$
Moreover, since the operators $\bm{U}_{p},K,$ and $l_{\bm{\gamma},\bm{\beta}}$ unitary, the set of functions $\{\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\}_{j=1}^{\infty}$ form an orthonormal basis for $L^{2}(G_{\bm{\beta}})$, and thus we can set
@@
From the above construction, the mean and covariance of the functions $\widetilde{h}_{j\bm{\beta}}$ under $G_{\bm{\beta}}$ are the same as those of $\widetilde{\phi}_{j\bm{\gamma}}$ under $F_{\bm{\gamma}}$. Thus, the empirical processes $v_{G,n}(\widetilde{h}_{j\bm{\beta}})$ and $v_{F,n}(\widetilde{\phi}_{j\bm{\gamma}})$ have the same standard limiting distribution.
The above results can be generalized to cases where the dimension of $\bm{\gamma}$ is smaller than $p$, say $q$. In this setting, one can simply expand the orthonormal set of score functions $\{a_{\bm{\gamma}_{k}}\}_{k=1}^{q}$ to a larger orthonormal set $\{a_{\bm{\gamma}_{k}}\}_{k=1}^{p}$ in $L^{2}(F_{\bm{\gamma}})$, ensuring that all elements remain orthogonal to the constant function $\bm{1}$. This extension can be accomplished, for instance, by selecting pqp-q additional functions from another orthonormal basis in $L^{2}(F_{\bm{\gamma}})$ outside the span of $\{\bm{1},a_{\bm{\gamma}_{1}},\dots,a_{\bm{\gamma}_{q}}\}$, and applying the Gram-Schmidt orthogonalization procedure.
While at first glance the above steps may appear rather burdensome, note that all operators involved in the K2 transform are linear. Moreover, to test the hypotheses in (2), test statistics based on the K2-orthonormal basis need to be evaluated only once on the observed data for each hypothesized model $G_{\bm{\beta}}$ being tested.
@@ 4.2 Simulation Studies
$0,10],\displaystyle q(x)=3u_{1}(x;\mu_{1},\sigma_{1})+5u_{2}(x;\mu_{2},\sigma_{2})+2u_{3}(x),\quad x\in\mathcal{X}=[-0$,0],
where $u_{1}$ and $u_{2}$ denote, respectively, the densities of truncated normal and truncated Laplace random variables; $u_{3}$ is the uniform density. The values of the location parameters are $\mu_{1}=-5$, $\mu_{2}=5$ and scale parameters are $\sigma_{1}=\sigma_{2}=3$. We aim to test the hypotheses in (2) for three different specifications of the null density, i.e.,
@@
$g_{\bm{\beta},3}(x)$ $=\beta_{1}u_{1}(x;-4,1)+\beta_{2}u_{2}(x;4,1)+(1-\beta_{1}-\beta_{2})u_{3}(x),$
where $\bm{\beta}=(\beta_{1},\beta_{2})$ is the unknown parameter vector to be estimated.
Their corresponding CDFs are denoted by $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, and $G_{\bm{\beta},3}$, respectively. The reference distribution, $F_{\bm{\gamma}}$, is chosen to be a truncated normal distribution over $\mathcal{X}$ with unknown parameter $\bm{\gamma}$ corresponding to its mean and variance. Figure 2 shows the histogram of the dataset considered, along with the densities of $q,g_{\bm{\gamma}},g_{\bm{\beta},1},g_{\bm{\beta},2},$ and $g_{\bm{\beta},3}$ estimated via maximum likelihood.
Figure 3: The simulated null distributions of the order selection statistic (left panels) and the subset selection test statistic (right panels), using basis functions obtained by composing the normalized shifted Legendre polynomials with the null CDFs (upper panels) and the K2 transform (lower panels), under $F_{\bm{\gamma}}$, $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, and $G_{\bm{\beta},3}$.
Consider the case in which the basis functions for $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, $G_{\bm{\beta},3}$ and $F_{\bm{\gamma}}$ are constructed as compositions of the normalized shifted Legendre polynomials on [0,1] with the null CDFs. Using these basis functions, we simulated the null distributions of the order selection and subset selection test statistics in (7) and (9) with $\widehat{S}_{m,n}$ chosen to be the unnormalized generalized score statistic in (6). Specifically, the null distributions of these statistics were simulated using the projected bootstrap described in (3.1), with 100,000 replicates and a maximum of $M_{n}=6$ basis functions. As shown in the upper panel of Figure 3, the null distributions of the order selection statistic under $G_{\bm{\beta},1}$, $G_{\bm{\beta},2},$ and $G_{\bm{\beta},3}$ differs significantly from that under $F_{\bm{\gamma}}$. While the distributions of the subset selection test statistic under $F_{\bm{\gamma}}$, $G_{\bm{\beta},1}$, and $G_{\bm{\beta},3}$ are rathe similar, they differ substantially from that obtained under $G_{\bm{\beta},2}$. These discrepancies are expected given that, in general, the unnormalized generalized score statistic in (6) is not distribution-free.
$H_{0}$ $\alpha=0.001$ $\alpha=0.05$ $\alpha=0.1$
@@
$F_{\bm{\gamma}}$ 0.645 0.315 0.939 0.826 0.967 0.902
$G_{\bm{\beta},1}$ 0.010 0.101 0.661 0.610 0.366 0.565 0.953 0.941 0.542 0.694 0.975 0.968
$G_{\bm{\beta},2}$ $2\times 10^{-4}$ 0.053 0.874 0.872 0.420 0.440 0.991 0.989 0.570 0.576 0.996 0.995
$G_{\bm{\beta},3}$ 0.005 0.026 0.037 0.026 0.159 0.266 0.349 0.302 0.277 0.383 0.469 0.427
Table 1: Comparing the power of the classical order selection and subset selection test statistics with their K2-transformed counterparts. The significance levels are $0.001,0.05,$ or $0.1$.
The same experiment was repeated, considering the K2 transformed basis functions for $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, $G_{\bm{\beta},3}$ constructed as described in Section 4.1. As shown in the lower panels of Figure 3, for both statistics considered, the simulated null distributions under $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, and $G_{\bm{\beta},3}$ are all indistinguishable from those under $F_{\bm{\gamma}}$. It follows that, for all models and statistics considered in this example, a sample size of 100 observations is sufficient to retrieve the distribution-free property even when two parameters are estimated.
Table 1 compares the power of test statistics constructed using the K2 basis functions with that obtained when considering compositions of normalized shifted Legendre polynomials with the null CDF. For all models under study, the K2-based test statistics exhibit higher power. However, this result should not be assumed to be true in general: the K2 transform leads to the construction of a new family of test statistics that can outperform the classical ones in certain settings but are not guaranteed to always lead to higher power.
@@ 5 Case study: analyzing an X-ray spectrum from RT Cru
where $\mathcal{X}=[1.65,2.05]$; $b(x)$ is a uniform background density on $\mathcal{X}$; and the functions $s_{r}(x)$ model each of the the three expected spectral lines. They consist of a convolution of a normal density with a Moffat function [^4] The Moffat function is equivalent to the density of a location-scale Student’s $t$-distribution with a location of 0, a scale of 0.025, and 4 degrees of freedom. (Moffat,, 1969) and specify as
@@
Figure 4: Left: $Q$ plots of the simulated order selection test statistics in (7) under $G_{\bm{\beta}}$ and in (18) under $F_{\bm{\gamma}}$. Right: $Q$ plots of the simulated subset selection test statistics in (9) under $G_{\bm{\beta}}$ and in (18) under $F_{\bm{\gamma}}$.
@@
where $p_{r}$ is the density of the truncated normal over $\mathcal{X}$, with $\mu_{r}$ being the known positions of the spectral lines and given after (25). The unknown parameter $\bm{\gamma}$ is estimated via the maximum likelihood. The basis functions for $F_{\bm{\gamma}}$ are chosen to be compositions of normalized shifted Legendre polynomials on $[0,1]$ with $F_{\bm{\gamma}}$. These functions are employed as a starting point to construct the K2 transformed basis functions $h_{j\bm{\beta}}$ for $G_{\bm{\beta}}$, whose residuals are subsequently used when calculating the order selection and subset selection statistics in (7) and (9), respectively, with $\widehat{S}_{m,n}$ as in (6). Their limiting null distributions are obtained by simulating 100,000 realizations of the statistics (18) under $F_{\bm{\gamma}}$ through the projected bootstrap (Cf. Section 3). The maximum number of basis functions used is $M_{n}=6$. To assess the accuracy of the approximation, we also simulated the null distributions of the statistics in (7) and (9) constructed using the K2 basis functions directly from $G_{\bm{\beta}}$. As shown by the QQ-plots in Figure 4, the resulting null distributions of (7) and (9) under $G_{\bm{\beta}}$ are indistinguishable from those of (18) under $F_{\bm{\gamma}}$.

### v3

@@ 4.1 On the construction of the K2 orthonormal basis
From Proposition 4, it follows that each function $\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$ fulfills (21), thus:
$\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\widetilde{\phi}_{j\bm{\gamma}}$ $=\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\left[\phi_{j\bm{\gamma}}-\bm{a}^{T}_{\bm{\gamma}}\left\langle\bm{a}_{\bm{\gamma}},\phi_{j\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}\right]$
$=\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}-\bm{b}^{T}_{\bm{\beta}}\left\langle\bm{b}_{\bm{\beta}},\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}}.$
We now have access to all the elements needed to define the K2 orthonormal basis at the core of the proposed distribution-freeness construction:
Proposition 5.
The set of functions $\{\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\}_{j=1}^{\infty}$ forms an orthonormal basis for $L^{2}(G_{\bm{\beta}})$. Define, for each $j\geq 1,$
@@
Then, the empirical processes $v_{\bm{\beta},n}(\widetilde{h}_{j\bm{\beta}})$ and $v_{\bm{\gamma},n}(\widetilde{\phi}_{j\bm{\gamma}})$ have the same limiting distribution under $G_{\bm{\beta}}$ and $F_{\bm{\gamma}}$, respectively.
Proof.
The first claim follows from the unitary nature of the operators $\bm{U}_{p}$, $K$, and $l_{\bm{\gamma},\bm{\beta}}$. The second claim is a direct consequence of the asymptotic Gaussianity of $v_{\bm{\beta},n}(\widetilde{h}_{j\bm{\beta}})$ and $v_{\bm{\gamma},n}(\widetilde{\phi}_{j\bm{\gamma}})$ and the construction of the $\{\widetilde{h}_{j\bm{\beta}}\}_{j=1}^{\infty}$ basis, which ensures that the empirical process indexed by such functions has, under $G_{\bm{\beta}}$, the same mean and covariance as the empirical process indexed by functions $\{\widetilde{\phi}_{j\bm{\gamma}}\}_{j=1}^{\infty}$ under $F_{\bm{\gamma}}$. ∎
The approach described above can be generalized to cases where the dimension of $\bm{\gamma}$ is smaller than $p$, say $s$. In this setting, one can simply expand the orthonormal set of score functions $\{a_{\bm{\gamma}_{k}}\}_{k=1}^{s}$ to a larger orthonormal set $\{a_{\bm{\gamma}_{k}}\}_{k=1}^{p}$ in $L^{2}(F_{\bm{\gamma}})$, ensuring that all elements remain orthogonal to each other and the constant function $\bm{1}$. This extension can be accomplished, for instance, by selecting psp-s additional functions from another orthonormal basis in $L^{2}(F_{\bm{\gamma}})$ outside the span of $\{\bm{1},a_{\bm{\gamma}_{1}},\dots,a_{\bm{\gamma}_{s}}\}$, and applying the Gram-Schmidt orthogonalization procedure.
@@ 4.2 Simulation Studies
$0,10],\displaystyle q(x)=3d_{1}(x;\mu_{1},\sigma_{1})+5d_{2}(x;\mu_{2},\sigma_{2})+2d_{3}(x),\quad x\in\mathcal{X}=[-0$,0],
where $d_{1}$ and $d_{2}$ denote, respectively, the densities of truncated normal and truncated Laplace random variables; $d_{3}$ is the uniform density. The values of the location parameters are $\mu_{1}=-5$, $\mu_{2}=5$ and scale parameters are $\sigma_{1}=\sigma_{2}=3$. We aim to test the hypotheses in (2) for three different specifications of the null density, i.e.,
@@
$g_{\bm{\beta},3}(x)$ $=\beta_{1}d_{1}(x;-4,1)+\beta_{2}d_{2}(x;4,1)+(1-\beta_{1}-\beta_{2})d_{3}(x),$
where $\bm{\beta}=(\beta_{1},\beta_{2})$ is the unknown parameter vector to be estimated. Their corresponding CDFs are denoted by $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, and $G_{\bm{\beta},3}$, respectively. The reference distribution, $F_{\bm{\gamma}}$, is chosen to be a truncated normal distribution over $\mathcal{X}$ with unknown parameter $\bm{\gamma}$ corresponding to its mean and variance. Figure 2 shows the histogram of the dataset considered, along with the densities of $q,g_{\bm{\gamma}},g_{\bm{\beta},1},g_{\bm{\beta},2},$ and $g_{\bm{\beta},3}$ estimated via maximum likelihood.
Figure 3: The simulated null distributions of the order selection statistics (left panels) and the subset selection test statistics (right panels), using basis functions obtained by composing the normalized shifted Legendre polynomials with the null CDFs (upper panels) and the K2 transform (lower panels), under $F_{\bm{\gamma}}$, $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, and $G_{\bm{\beta},3}$.
Consider the case in which the basis functions used in (1) are constructed as compositions of the normalized shifted Legendre polynomials on [0,1] and the CDF under the null, i.e.,
$\begin{split}h_{1,\beta}(x)&=\sqrt{12}[G_{\bm{\beta},s}(x)-0.5],\\ h_{2,\beta}(x)&=\sqrt{5}[6G^{2}_{\bm{\beta},s}(x)-6G_{\bm{\beta},s}(x)+1],\\ h_{3,\beta}(x)&=\sqrt{7}[20G^{3}_{\bm{\beta},s}(x)-30G^{2}_{\bm{\beta},s}(x)+12G_{\bm{\beta},s}(x)-1],\\ h_{4,\beta}(x)&=3[70G^{4}_{\bm{\beta},s}(x)-140G^{3}_{\bm{\beta},s}+90G^{2}_{\bm{\beta},s}(x)-20G_{\bm{\beta},s}(x)+1],\quad\text{etc.}\end{split}$ (24)
with $s=1,2,3$ when testing $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, and $G_{\bm{\beta},3}$, respectively. For each of the three null models considered, we simulated the null distributions of the statistics $\widehat{T}_{n}$ in (7) and $\widetilde{T}_{n}$ in (9) constructed using the above-mentioned sets of basis functions, up to $M_{n}=6$, and with $\widehat{S}_{m,n}$ chosen to be the unnormalized generalized score statistic in (6). The simulation was performed using the projected bootstrap, described in Section 3.1, with 100,000 replicates. For comparison, the distribution of the statistics in (18) under $F_{\bm{\gamma}}$ has also been simulated following the same sampling scheme and choosing the basis functions $\phi_{j}$ to be a composition of normalized shifted Legendre polynomials on $[0,1]$ and $F_{\bm{\gamma}}$; hence, for $j=1,\dots,4$ they correspond to the right-hand sides of the equations in (24) with $F_{\bm{\gamma}}$ in place of $G_{\bm{\beta},s}$. As shown in the upper panel of Figure 3, the null distributions of the order selection statistic $\widehat{T}_{n}$ under $G_{\bm{\beta},1}$, $G_{\bm{\beta},2},$ and $G_{\bm{\beta},3}$ differs substantially from that constructed under $F_{\bm{\gamma}}$, and given by the first expression in (18). In the case of the subset selection test statistic $\widetilde{T}_{n}$, the null distributions under $G_{\bm{\beta},1}$, and $G_{\bm{\beta},3}$ are rather similar. They also match the null distribution of the same statistic, constructed under $F_{\bm{\gamma}}$, and given by the right expression in (18). However, all these distributions differ substantially from that of $\widetilde{T}_{n}$ under $G_{\bm{\beta},2}$. These discrepancies are expected given that, in general, the unnormalized generalized score statistic in (6) is not distribution-free.
$\alpha=0.001$ $\alpha=0.05$ $\alpha=0.1$
@@
$F_{\bm{\gamma}}$ 0.859 0.851 1.000 1.000 1.000 1.000
$G_{\bm{\beta},1}$ 0.008 0.091 0.053 0.025 0.351 0.551 0.519 0.333 0.533 0.688 0.662 0.473
$G_{\bm{\beta},2}$ 0.016 0.038 0.643 0.695 0.238 0.424 0.967 0.969 0.357 0.566 0.985 0.986
$G_{\bm{\beta},3}$ 0.007 0.028 0.037 0.035 0.164 0.285 0.371 0.351 0.286 0.402 0.498 0.482
Table 1: Power comparison between classical order selection and subset selection test statistics with their K2-transformed counterparts. The null distributions are $G_{\bm{\beta},1},G_{\bm{\beta},2},G_{\bm{\beta},3}$; the reference distribution is $F_{\bm{\gamma}}$; the true data-generating distribution is $Q$; and significance levels are set at $0.001,0.05,$ and $0.1$.
To assess the impact of the K2 transform, we repeated the same experiment but for the case in which the basis functions used in (1) when testing $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, $G_{\bm{\beta},3}$ are the K2-transformed basis functions, obtained as described in Section 4.1, and corresponding to the images of the K2 transform applied to the basis functions $\phi_{j}$ given by compositions of Legendre polynomials and $F_{\bm{\gamma}}$. In what follows, we denote the order selection and subset selection statistics constructed using the K2-transformed basis functions by $\widehat{T}^{K}_{n}$ and $\widetilde{T}^{K}_{n}$, respectively. As shown in the lower panels of Figure 3, for both such statistics, the simulated null distributions under $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, and $G_{\bm{\beta},3}$ are indistinguishable from the distribution of the corresponding statistics in (18) simulated under $F_{\bm{\gamma}}$. It follows that, for all models and statistics considered in this example, a sample size of 100 observations is sufficient to retrieve the distribution-free property even when two parameters are estimated.
To evaluate the performance of the testing procedure in (2) under different choices of basis functions, we computed the simulated power using both order selection and subset selection statistics and calculated it as the proportion of bootstrap replications under $Q$ for which the value of these statistics exceeds the critical value obtained from the simulated null distribution at the prescribed significance levels. Table 1 presents the resulting power when testing $G_{\bm{\beta},1},G_{\bm{\beta},2},G_{\bm{\beta},3}$ using the statistics $\widehat{T}_{n}$, $\widetilde{T}_{n}$, $\widehat{T}^{K}_{n}$, and $\widetilde{T}^{K}_{n}$ at significance levels $\alpha=0.001,0.05,$ and $0.1$. The power when testing the reference distribution $F_{\bm{\gamma}}$ using $\widehat{T}_{n}$ and $\widetilde{T}_{n}$ is also reported. When testing $G_{\bm{\beta}_{2}}$ and $G_{\bm{\beta}_{3}}$, the K2-based test statistics $\widehat{T}^{K}_{n}$ and $\widetilde{T}^{K}_{n}$ exhibit substantially higher power than $\widehat{T}_{n}$ and $\widetilde{T}_{n}$. Notably, when testing $G_{\bm{\beta},2}$ at a significance level of $\alpha=0.001$, $\widetilde{T}_{n}^{K}$ attains power as high as 0.695, whereas $\widetilde{T}_{n}$ achieves only 0.038. When testing $G_{\bm{\beta}_{1}}$, $\widehat{T}^{K}_{n}$ shows higher power than $\widehat{T}_{n}$ while the power of $\widetilde{T}^{K}_{n}$ is lower than that of $\widetilde{T}_{n}$. This reflects the fact that K2-transformed statistics can exhibit higher power against certain alternatives, though such improvements in power should not be assumed to be true in general.
@@ 5 Case study: analyzing an X-ray spectrum from RT Cru
where $\mathcal{X}=[1.65,2.05]$; $b(x)$ is a uniform background density on $\mathcal{X}$; and the functions $s_{r}(x)$ model each of the the three expected spectral lines. They consist of a convolution of a normal density with a Moffat function [^5] The Moffat function is equivalent to the density of a location-scale Student’s $t$-distribution with a location of 0, a scale of 0.025, and 4 degrees of freedom. (Moffat, 1969) and specify as
@@
Figure 4: Left: $Q$ plots of the simulated order selection test statistics in (7) under $G_{\bm{\beta}}$ and in (18) under $F_{\bm{\gamma}}$. Right: $Q$ plots of the simulated subset selection test statistics in (9) under $G_{\bm{\beta}}$ and in (18) under $F_{\bm{\gamma}}$. For both statistics computed under $G_{\bm{\beta}}$, the functions $h_{j\bm{\beta}}$ are elements of the $K2$ basis costructed as in Proposition 5.
@@
where $p_{r}$ is the density of a truncated normal over $\mathcal{X}$, with $\mu_{r}$ being the known positions of the spectral lines and listed after (26). The unknown parameter $\bm{\gamma}=(\gamma_{1},\gamma_{2},\gamma_{3})$ is estimated via the maximum likelihood. The basis functions for $F_{\bm{\gamma}}$ are chosen to be compositions of normalized shifted Legendre polynomials on $[0,1]$ with $F_{\bm{\gamma}}$. These functions are employed as a starting point to construct the K2-transformed basis functions $h_{j\bm{\beta}}$ for $L^{2}(G_{\bm{\beta}})$, whose residuals are subsequently used when calculating the order selection and subset selection statistics in (7) and (9), respectively, with $\widehat{S}_{m,n}$ as in (6) and $\{h_{j\bm{\beta}}\}_{j=1}^{M_{n}}$ as in Proposition 5. Hence, consistently with the notation used in the previous section, we denote them with $\widehat{T}_{n}^{K}$ and $\widetilde{T}_{n}^{K}$. Their limiting null distributions are obtained by simulating 100,000 realizations of the statistics (18) under $F_{\bm{\gamma}}$ through the projected bootstrap (cf. Section 3). The maximum number of basis functions used is $M_{n}=6$. To assess the accuracy of the approximation, we also simulated the distributions of $\widehat{T}_{n}^{K}$ and $\widetilde{T}_{n}^{K}$ under $G_{\bm{\beta}}$. As shown by the QQ-plots in Figure 4, the null distributions of such statistics under $G_{\bm{\beta}}$ are indistinguishable from those of the statistics in (18) under $F_{\bm{\gamma}}$.